OpenStudy (anonymous):

a 15 ft ladder is leaning against a building. If the bottom of the ladder is 9 ft from the base of the building, how high does the ladder reach

OpenStudy (anonymous):

this is solvable by pythagoras' theorem, in this case the hypotenuse of rt angld triangle is the ladder so 15^2 = 9^2 + h^2 where h is what you require

OpenStudy (anonymous):

The answer is 12 ft but I dont know how the instructor got that

OpenStudy (anonymous):

Use the pythagorean theorem \[15^{2} = h ^{2}+9 ^{2}\] \[h = \sqrt{15^{2}-9^{2}} = 12\]

OpenStudy (anonymous):

thank you